OpenStudy (anonymous):

What is 900 feet per day converted into yards per hour?

OpenStudy (anonymous):

Multiply by (1 yard/3 feet) and multiply again by (1 day/24 hours). This will do a couple of things. It will cancel and introduce the appropriate units of measure. And this is quite allowable because you are in essence just multiplying by "1" in each case because the above 2 factors are identities and are each essentially "1".

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @12rabbits

OpenStudy (anonymous):

thx & ur welcome

OpenStudy (tkhunny):

\(\dfrac{ft}{day}\cdot\dfrac{1\;yd}{3\;ft}\cdot\dfrac{1\;day}{24\;hr}\)